The Adams Management Association held its annual public relations luncheon in April 2017. Based on the previous year’s results, the organization allocated $21,842 of its operating budget to cover the cost of the luncheon. To ensure that costs would be appropriately controlled, Molly Hubbard, the treasurer, prepared the following budget for the 2017 luncheon.

The budget for the luncheon was based on the following expectations.

  1. The meal cost per person was expected to be $11.90. The cost driver for meals was attendance, which was expected to be 1,410 individuals.
  2. Postage was based on $0.46 per invitation and 3,050 invitations were expected to be mailed. The cost driver for postage was number of invitations mailed.
  3. The facility charge is $1,100 for a room that will accommodate up to 1,600 people; the charge for one to hold more than 1,600 people is $1,600.
  4. A fixed amount was designated for printing, decorations, the speaker’s gift, and publicity.
ADAMS MANAGEMENT ASSOCIATION
Public Relations Luncheon Budget
April 2017
Operating funds allocated $ 21,842
Expenses
Variable costs
Meals (1,410 × $11.90) 16,779
Postage (3,050 × 0.46) 1,403
Fixed costs
Facility 1,100
Printing 960
Decorations 850
Speaker's gift 140
Publicity 610
Total expenses 21,842
Budget surplus (deficit) $ 0

Actual results for the luncheon follow.

ADAMS MANAGEMENT ASSOCIATION
Actual Results for Public Relations Luncheon
April 2017
Operating funds allocated $ 21,842
Expenses
Variable costs
Meals (1,630 × $12.60) 20,538
Postage (4,050 × 0.46) 1,863
Fixed costs
Facility 1,600
Printing 960
Decorations 850
Speaker's gift 140
Publicity 610
Total expenses 26,561
Budget deficit $ (4,719 )

Reasons for the differences between the budgeted and actual data follow.

  1. The president of the organization, Zachary Taylor, increased the invitation list to include 1,000 former members. As a result, 4,050 invitations were mailed.
  2. Attendance was 1,630 individuals. Because of higher-than-expected attendance, the luncheon was moved to a larger room, thereby increasing the facility charge to$1,600
  3. At the last minute, Ms. Cole decided to add a dessert to the menu, which increased the meal cost to $12.6 per person.
  4. Printing, decorations, the speaker’s gift, and publicity costs were as budgeted.

Required:

a. Prepare a flexible budget and compute the sales and variable cost volume variances based on a comparison between the master budget and the flexible budget.

b. Compute flexible budget variances by comparing the flexible budget with the actual results.
Answer #1
A.Flexible Budget:
Operating funds allocated 21842
Expenses
Variable costs
Meals (1630*11.90) 19397
Postage (4050*0.46) 1863
Fixed costs
Facility 1600
Printing 960
Decorations 850
Speaker’s gift 140
Publicity 610
Total expenses 4160
Budget surplus (deficit) -3578
Sales volume variance=21842-21842=0
Variable cost as per Master budget=16779+1403=18182
Variable cost as per Flexible budget=19397+1863=21260
Variable cost volume variance=21260-18182=3078 U
B.Comparison of Flexible with Actual Flexible Actual Variance
Budget
Operating funds allocated 21842 21842 0
Expenses
Variable costs
Meals (1630*11.9) 19397 20538 1141 U
Postage (4050*0.46) 1863 1863 0
Fixed costs
Facility 1600 1600 0
Printing 960 960 0
Decorations 850 850 0
Speaker’s gift 140 140 0
Publicity 610 610 0
Total expenses 4160 4160 1141 U
Budget surplus (deficit) -3578 -4719 1141 U